# Deconstructing the Goldilocks Zone of Neural Network Initialization

### v1

@@ 4 Features of the Goldilocks Zone
In this section, we inquire about the properties of the interior of the Goldilocks zone. In particular, we will associate extreme values of positive curvature with certain features of the network, initialization, and the data. To this end, we study the spectral properties of the G-term that, provided Equation 6 holds, transfer to the loss Hessian as well. Equation 5 reveals that $G^{\prime}_{*}(p)\propto G_{*}(p)$ for the same $p$, so that structural changes in the eigenspectrum of the G-term at scaled initializations are due to the changing softmax output $p$ alone. Therefore, we can drop the superscript and study the behavior of $G_{*}$ as a function of $p$, which is the goal of this section. In particular, we will find that high positive curvature is associated with low model confidence, which, in turn, leads to vanishing expected loss gradients for balanced datasets and low initial loss.
@@
At the basis of our analysis is the random model introduced by Fort & Ganguli (2019). They observed that same-logit gradients cluster across training samples, and that the $K$ corresponding logit gradient means $c_{k}$ are nearly orthogonal to each other. In light of this observation, they model $c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{P})$ and, for every input $X^{\mu}$, let the corresponding $k$-th logit gradient be $\nabla_{\theta}f_{k}(X^{\mu})=c_{k}+e^{\mu}_{k}$ with iid residuals $e_{k}^{\mu}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{P})$. Recall from Section 2 that, in practice, we compute positive curvature of the Hessian projected to a low-rank hyperplane given by an orthonormal basis $R\in\mathbb{R}^{P\times d}$, so we actually care about gradients with respect to the $d$-dimensional trainable parameters. Conveniently, the random model still holds in this low-rank subspace since the corresponding gradients are just $R^{\top}c_{k}+R^{\top}e_{k}^{\mu}$ where $R^{\top}c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{d})$ and $R^{\top}e^{\mu}_{k}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{d})$ as linear transformations of isotropic Gaussian variables by $R^{\top}$ satisfying $R^{\top}R=I_{d}$. Thus, we drop the transformation and simply assume $c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{d})$ and $e^{\mu}_{k}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{d})$. Note that, in particular, the variance of logit gradients and the corresponding residuals does not depend on $d$. We make two more simplifying assumptions and require $c_{k}$ to be pairwise orthogonal and have equal squared length $\lVert c_{k}\rVert^{2}=\mathbb{E}\lVert c_{k}\rVert^{2}=d\sigma^{2}_{c}$.
@@
Figure 3: Positive curvature of the true G-term $G_{*}$ and the expected G-term $\mathbb{E}G_{*}$ from Equation 10 computed using logits associated with different initialization scales (model confidence). We used a low-rank subspace with $d=50$. Error bands represent min/max across $3$ seeds. Left: LeNet-300-100 on FashionMNIST; Right: LeNet-5 on CIFAR-10.
@@
For a single training sample, let $C$ and $E$ be $k\times d$ matrices of logit gradient means and the corresponding residuals (i.e., $C_{k}=c_{k}$ and $E_{k}=e_{k}$), respectively; the Jacobian is then $J=C+E$ and so $G_{*}=(C+E)^{\top}[\text{diag}(p)-pp^{\top}](C+E)$. Taking expectation over data, we get
$\mathbb{E}_{e}G_{*}$ $=C^{\top}[\text{diag}(p)-pp^{\top}]C+(1-\lVert p\rVert^{2})\sigma_{e}^{2}I_{d}$
@@
where in the second line, we switch to the orthonormal basis of the normalized logit gradient means $c_{k}$. We can now compute the positive curvature of $\mathbb{E}_{e}G_{*}$ directly from its elements. Assuming that $p$ is not one-hot, we get
$\frac{\text{Tr}(\mathbb{E}_{e}G_{*})}{\lVert\mathbb{E}_{e}G_{*}\rVert_{F}}=\frac{\sqrt{d}(\sigma_{c}^{2}+\sigma_{e}^{2})}{\sqrt{\sigma_{e}^{4}+2\sigma_{e}^{2}\sigma_{c}^{2}+d\sigma_{c}^{4}{\Gamma}_{p}^{-2}}}$ (10)
where $\Gamma_{p}$ is the positive curvature of $\text{diag}(p)-pp^{\top}$ as defined in Equation 2. For the expected G-term computed on a batch of samples with potentially different softmax probabilities $p^{\mu}$, $\Gamma_{p}$ computes positive curvature of the average matrix $\frac{1}{B}\sum_{\mu=1}^{B}\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$. Having empirically estimated values $\sigma_{c}^{2}$ and $\sigma_{e}^{2}$, we use Equation 10 to compute the positive curvature of $\mathbb{E}_{e}G_{*}$ on real softmax outputs $p$ found by scaling initialization, and find that it approximates the positive curvature of the true G-term across varying model confidence levels very well (Figure 3).
Figure 4: Left: The dependence of positive curvature on the dimension $d$ of the low-rank hyperplane used to compute the projected G-term. Right: $\Gamma_{p}^{-2}$ of the matrix $\text{diag}(p)-pp^{\top}$ where vectors $p$ are produced by scaling $30$ different logit sets by $\alpha\in[10^{-2},10^{2}]$ ($30$ pink curves). The red curve corresponds to $\Gamma_{p}^{-2}$ computed for the average matrix $\frac{1}{30}\sum_{\mu=1}^{30}\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$.
@@
We can now explain the dependence of positive curvature on the dimension $d$ of the hyperplane used to project the model parameters, which was first noticed by Fort & Scherlis (2019), see the left plot in Figure 4. Indeed, according to Equation 10, larger $d$ are associated with higher positive curvature of the expected G-term given the same softmax outputs $p$.
@@
Figure 2 shows that positive curvature of the G-term monotonically decreases, reaching zero at some $\alpha>1$ due to a collapsing softmax. Since larger $\alpha$ are associated with colder softmax distributions, it is natural to hypothesize that positive curvature of the G-term is inversely related to model confidence. Equation 10 reveals that positive curvature of the expected G-term is directly related to that of $\text{diag}(p)-pp^{\top}$ and is larger for smaller values of $\Gamma_{p}^{-2}$. Despite the seemingly nice form of this matrix, $\Gamma_{p}$ is difficult to analyze algebraically; still, a few general remarks shed light on the above relationship. First, Section 4 can be used to show that $\mathbb{E}G_{*}$ has a “bi-level” eigenspectrum that yields maximal positive curvature when the softmax output $p$ is uniform (Appendix B). Second, as seen from Equation 2, positive curvature of any rank-$r$ positive semidefinite matrix is just the ratio of L1 and L2 norms of its $r$ non-zero eigenvalues, which is known to fall between $1$ and $\sqrt{r}$. Therefore, $\Gamma_{p}\in[1,\sqrt{K-1}]$, and it achieves its maximal value when $p$ is uniform. In Appendix C, we prove that $\Gamma_{p}\rightarrow 1$ as $p$ collapses to a one-hot vector due to the increasing initialization scale $\alpha$. Moreover, Figure 4 (right) reveals that $\Gamma_{p}$ is, in fact, monotonic between its extreme values with respect to the entropy of $p$. Overall, this suggests that $\Gamma^{-2}_{p}$ in Equation 10 grows with model confidence, which supports our hypothesis.
Figure 5: The effects of the average softmax output $\hat{Q}$ and the actual target prior $Q$ on the batch gradient. Left: Cosine similarity of gradients computed on different datasets (SVHN, CIFAR-10, and a randomly generated dataset) of an unscaled LeNet-5. Right: Given a fixed unscaled initialization of LeNet-5, we sample $2,000$ different class priors $Q$ uniformly from a probability simplex $\Delta_{10}$. For each of them, we choose a subset of CIFAR-10 with $5,000$, compute the cross-entropy gradient, and plot its norm against $\lVert\hat{Q}-Q\rVert_{2}$. As we predicted, the norms follow a linear trend with the predicted slope ($\sigma_{c}^{2}$ was estimated on the entire CIFAR-10 dataset; $d=P=61,170$.
@@
In Section 3, we saw that vanishing logit gradients associated with small-norm network initializations diminish the role of the G-term in the Gauss-Newton decomposition, leading to zero positive curvature of the Hessian. The term “vanishing gradients” commonly refers to the condition of signal decay during backpropagation to deeper layers, emerging either due to saturated activation functions or small spectral radii of parameter matrices (Bengio et al., 1994; Pascanu et al., 2013). The vanishing logit gradients observed at low initialization scales of homogeneous networks fall in the latter category. Using the random model of Fort & Ganguli (2019), we derive yet another, previously unknown type of vanishing cross-entropy gradients caused by the match between the average softmax output of the model $\hat{Q}=\mathbb{E}p\approx\frac{1}{N}\sum_{\mu}p^{\mu}$ and the dataset prior $Q=\mathbb{E}y\approx\frac{1}{N}\sum_{\mu}y^{\mu}$ where expectation is taken over the data distribution. To this end, consider the expected batch cross-entropy loss gradient:
$\mathbb{E}_{E}\left[\frac{\partial\ell}{\partial\theta}\right]$ $=\mathbb{E}_{E}\left[\frac{1}{B}\sum_{\mu=1}^{B}\sum_{k=1}^{K}(c_{k}+e_{k}^{\mu})(p_{k}^{\mu}-y_{k}^{\mu})\right]$
@@
By the orthogonality assumption on logit gradient means $c_{k}$, the norm of this quantity is $d\sigma_{c}^{2}\lVert\hat{Q}-Q\rVert$. In Figure 5 (right), we confirm this linear relationship on LeNet-5 shockingly well. Normally, randomly-initialized models are biased and have non-uniform average priors $\hat{Q}$, so that $\lVert\hat{Q}-Q\rVert>0$. In this case, we expect loss gradients computed on two potentially unrelated but balanced datasets to have considerable overlap as both should align with the non-zero expected gradient. Indeed, Figure 5 (left) shows that gradients of LeNet-5 computed on CIFAR-10 and on random images drawn from a Gaussian have cosine similarity $>0.7$, which is quite noteworthy for a 61,170-dimensional space. In constrast, for low-confident models that always predict uniform distribution (e.g., when $\alpha\ll 1$), logit gradients cancel each other out when computed on a large balanced batch. Hence, rather counterintuitively, we derive that high positive curvature is associated with low-norm cross-entropy gradients for balanced datasets, as we verified in Figure 6.
@@ 5 Connections to Optimization
We optimize $\alpha$-scaled homogeneous networks using vanilla full-batch gradient descent for 20,000 epochs. In particular, we use a fully-connected LeNet-300-100 with two hidden layers on FashionMNIST and a convolutional LeNet-5 with $4$ hidden layers on CIFAR-10 (LeCun et al., 1998), all implemented in PyTorch (Paszke et al., 2017). We set softmax temperature to $T=1$ to link the Goldilocks zone to the initialization norm. Recall from Section 3 that logit gradients of the $\alpha$-scaled network $f^{\prime}$ are $\alpha^{L-1}$ times the respective logit gradients of the unscaled model $f$, so we multiply a preset learning rate $\eta_{0}$ by $\alpha^{2-L}$ to ensure that updates are initially commensurate to the weights of $f^{\prime}$. In fact, this correction factor ensures that $f^{\prime}$ has exactly the same training dynamics as $f$ provided that logit gradients are combined using the same softmax output $p=p^{\prime}$ to compute the update vector as indicated in Equation 3. From this perspective, our experiments essentially constitute an ablation study examining the effects of extreme model (un)confidence, which corresponds to the exterior of the Goldilocks zone in our setup.
@@
Our findings are summarized in Figure 7. As long as the softmax ouptut remains uniform ($\alpha<1$), homogeneous models remain in the divergence regime characterized by an increase of the parameter norm in accordance with previous studies (Liu et al., 2023). Once the weights become sufficiently large for the predictions to become non-uniform, the network circles back to the Goldilocks zone and trains normally thereafter provided that the learning rate is approximately admissible, i.e., $\eta=\eta_{0}\alpha^{2-L}<\lVert H\rVert_{2}$ (Lewkowycz et al., 2020; Cohen et al., 2021; Wang et al., 2022). Hence, in Figure 8 (left) we observe a linear separation between trainable and divergent setups that extends well into the low-norm initialization region and way beyond the Goldilocks zone (see Appendix E for further discussion). When the learning rate is too high, however, the network breezes past the Goldilocks zone and either diverges to infinity or develops an increasing amount of zero logits (middle column in Figure 8). This happens as some negative weights grow much faster than the positive ones, promoting zero activations within the network by virtue of ReLU.
@@
The observations above reveal that homogeneous networks may not only converge when initialized outside the Goldilocks zone but also fail to learn when optimized well within its boundaries (Figure 8). The former occurs as the model circles back to the Goldilocks zone ($\alpha\ll 1$) or when it learns the lazy solution ($\alpha\gg 1$), and the latter—as it develops zero logits. Indeed, LeNet-5 does not perform above random for any $\alpha>1.8$ and any learning rate, whereas the corresponding Goldilocks zone spans $\alpha\in[1,18]$. Lastly, Figure 8 simply illustrates that phase transitions between different optimization regimes are not necessarily aligned with the Goldilocks zone, which is naturally highlighted in the right plots. Thus, we argue that positive curvature of the loss Hessian is not an appropriate metric to evaluate the quality of initializations in terms of network trainability and generalization.

### v2

@@ 4 Features of the Goldilocks Zone
In this section, we inquire about the properties of the interior of the Goldilocks zone. In particular, we will associate extreme values of positive curvature with certain features of the network, initialization, and the data. To this end, we study the spectral properties of the G-term that, provided Equation 6 holds, transfer to the loss Hessian as well. Equation 5 reveals that $\mathcal{G}^{\prime}(p)\propto\mathcal{G}(p)$ for the same $p$, so that structural changes in the eigenspectrum of the G-term at scaled initializations are due to the changing softmax output $p$ alone. Therefore, we can drop the superscript and study the behavior of $\mathcal{G}$ as a function of $p$, which is the goal of this section. In particular, we will find that positive curvature is associated with low model confidence, which, in turn, leads to vanishing expected loss gradients for balanced datasets and low initial loss.
@@
At the basis of our analysis is the random model introduced by Fort & Ganguli (2019). They observed that same-logit gradients cluster across training samples, and that the $K$ corresponding logit gradient means $c_{k}$ are nearly orthogonal to each other. In light of this observation, they model $c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{P})$ and, for every input $X^{\mu}$, let the corresponding $k$-th logit gradient be $\nabla_{\theta}f_{k}(X^{\mu})=c_{k}+e^{\mu}_{k}$ with iid residuals $e_{k}^{\mu}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{P})$. Recall from Section 2 that, in practice, we compute positive curvature by projecting the Hessian onto a low-rank hyperplane given by an orthonormal basis $R\in\mathbb{R}^{P\times d}$, so we actually care about gradients with respect to the $d$-dimensional trainable parameters. Conveniently, the random model still holds in this low-rank subspace since the corresponding gradients are just $R^{\top}c_{k}+R^{\top}e_{k}^{\mu}$ where $R^{\top}c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{d})$ and $R^{\top}e^{\mu}_{k}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{d})$ as linear transformations of isotropic Gaussian variables by $R^{\top}$ satisfying $R^{\top}R=I_{d}$. Thus, we drop the transformation and simply assume $c_{k}\sim\mathcal{N}(0,\sigma_{c}^{2}I_{d})$ and $e^{\mu}_{k}\sim\mathcal{N}(0,\sigma_{e}^{2}I_{d})$. Note that, in particular, the variance of logit gradients and the corresponding residuals does not depend on $d$. We make two more simplifying assumptions and require $c_{k}$ to be pairwise orthogonal and have equal squared length $\lVert c_{k}\rVert^{2}=\mathbb{E}\lVert c_{k}\rVert^{2}=d\sigma^{2}_{c}$.
@@
Figure 3: Excess of positive eigenvalues of the true G-term $\mathcal{G}$ and the expected G-term $\mathbb{E}\mathcal{G}$ from Equation 8 computed using logits associated with different initialization scales (model confidence). We used a low-rank subspace with $d=50$. Error bands represent min/max across $3$ seeds. Left: LeNet-300-100 on FashionMNIST; Right: LeNet-5 on CIFAR-10.
@@
For a single training sample, let $C$ and $E$ be $k\times d$ matrices of logit gradient means and the corresponding residuals (i.e., $C_{k}=c_{k}$ and $E_{k}=e_{k}$), respectively; the Jacobian is then $J=C+E$ and so $\mathcal{G}=(C+E)^{\top}[\text{diag}(p)-pp^{\top}](C+E)$. Taking expectation over data, we get
$\mathbb{E}_{e}\mathcal{G}$ $=C^{\top}[\text{diag}(p)-pp^{\top}]C+(1-\lVert p\rVert^{2})\sigma_{e}^{2}I_{d}$
@@
Figure 4: Left: The dependence of excess of positive eigenvalues on the dimension $d$ of the low-rank hyperplane used to compute the projected G-term. Right: $\Gamma_{p}^{-2}$ of the matrix $\text{diag}(p)-pp^{\top}$ where vectors $p$ are produced by scaling $30$ different logit sets by $\alpha\in[10^{-2},10^{2}]$ ($30$ pink curves). The red curve corresponds to $\Gamma_{p}^{-2}$ computed for the average matrix $\frac{1}{30}\sum_{\mu=1}^{30}\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$.
Figure 5: The effects of the average softmax output $\hat{Q}$ and the actual target prior $Q$ on the batch gradient. (a): Cosine similarity of gradients computed on different datasets (SVHN, CIFAR-10, and a randomly generated dataset) at the same initialization of LeNet-5 (downscaled by $\alpha=0.01$ when $\hat{Q}=\text{uniform}$). To achieve $Q\neq\text{uniform}$, we inject artificial class imbalance by subsampling datasets using a procedure suggested by Cui et al. (2019). (b): Given a fixed unscaled initialization of LeNet-5, we sample $2,000$ different class priors $Q$ uniformly from a probability simplex $\Delta_{10}$. For each of them, we sample a subset of CIFAR-10 with $5,000$, compute the cross-entropy gradient, and plot its norm against $\lVert\hat{Q}-Q\rVert_{2}$. As we predicted, the norms follow a linear trend with the predicted slope ($\sigma_{c}^{2}$ was estimated on the entire CIFAR-10 dataset; $d=P=61,170$).
where in the second line, we switch to the orthonormal basis of the normalized logit gradient means $c_{k}$. We can now compute the excess of positive eigenvalues of $\mathbb{E}_{e}\mathcal{G}$ directly from its elements. Assuming that $p$ is not one-hot, we get
$\frac{\text{Tr}(\mathbb{E}_{e}\mathcal{G})}{\lVert\mathbb{E}_{e}\mathcal{G}\rVert_{F}}=\frac{\sqrt{d}(\sigma_{c}^{2}+\sigma_{e}^{2})}{\sqrt{\sigma_{e}^{4}+2\sigma_{e}^{2}\sigma_{c}^{2}+d\sigma_{c}^{4}{\Gamma}_{p}^{-2}}}$ (8)
where $\Gamma_{p}$ is the excess of positive eigenvalues of $\text{diag}(p)-pp^{\top}$ as defined in Equation 2. For the expected G-term computed on a batch of samples with potentially different softmax probabilities $p^{\mu}$, $\Gamma_{p}$ computes excess of positive eigenvalues of the average matrix $\frac{1}{B}\sum_{\mu=1}^{B}\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$. Having empirically estimated values $\sigma_{c}^{2}$ and $\sigma_{e}^{2}$, we use Equation 8 to compute the excess of positive eigenvalues of $\mathbb{E}_{e}\mathcal{G}$ on real softmax outputs $p$ found by scaling initialization, and find that it approximates the excess of positive eigenvalues of the true G-term across varying model confidence levels very well (Figure 3).
@@
We can now explain the dependence of positive curvature on the dimension $d$ of the hyperplane used to project the model parameters, which was first noticed by Fort & Scherlis (2019), see the left plot in Figure 4. Indeed, according to Equation 8, larger $d$ are associated with higher excess of positive eigenvalues of the expected G-term given the same softmax outputs $p$.
@@
Figure 2 shows that excess of positive eigenvalues of the G-term monotonically decreases, reaching zero at some $\alpha>1$ due to a collapsing softmax. Since larger $\alpha$ are associated with colder softmax distributions, it is natural to hypothesize that excess of positive eigenvalues of the G-term is inversely related to model confidence. Equation 8 reveals that excess of positive eigenvalues of the expected G-term is directly related to that of $\text{diag}(p)-pp^{\top}$ and is larger for smaller values of $\Gamma_{p}^{-2}$. Despite the seemingly nice form of this matrix, $\Gamma_{p}$ is difficult to analyze algebraically; still, a few general remarks shed light on the above relationship. First, Section 4 can be used to show that $\mathbb{E}\mathcal{G}$ has a “bi-level” eigenspectrum that yields maximal excess of positive eigenvalues when the softmax output $p$ is uniform (Appendix B). Second, as seen from Equation 2, excess of positive eigenvalues of any rank-$r$ positive semidefinite matrix is just the ratio of L1 and L2 norms of its $r$ non-zero eigenvalues, which is known to fall between $1$ and $\sqrt{r}$. Therefore, $\Gamma_{p}\in[1,\sqrt{K-1}]$, and it achieves its maximal value when $p$ is uniform. In Appendix C, we prove that $\Gamma_{p}\rightarrow 1$ as $p$ collapses to a one-hot vector due to the increasing initialization scale $\alpha$. Moreover, Figure 4 (right) reveals that $\Gamma_{p}$ is, in fact, monotonic between its extreme values with respect to the entropy of $p$. Overall, this suggests that $\Gamma^{-2}_{p}$ in Equation 8 grows with model confidence, which supports our hypothesis.
@@
In Section 3, we saw that vanishing logit gradients associated with small-norm network initializations diminish the role of the G-term in the Gauss-Newton decomposition, leading to zero positive curvature of the loss. The term “vanishing gradients” commonly refers to the condition of signal decay during backpropagation to deeper layers, emerging either due to saturated activation functions or small spectral radii of parameter matrices (Bengio et al., 1994; Pascanu et al., 2013). The vanishing logit gradients observed at low initialization scales of homogeneous networks fall in the latter category. Using the random model of Fort & Ganguli (2019), we derive yet another, previously unknown type of vanishing cross-entropy gradients caused by the match between the average softmax output of the model $\hat{Q}=\mathbb{E}p\approx\frac{1}{N}\sum_{\mu}p^{\mu}$ and the dataset prior $Q=\mathbb{E}\text{OH}(y)\approx\frac{1}{N}\sum_{\mu}\text{OH}(y^{\mu})$ where expectation is taken over the data distribution. To this end, consider the expected batch cross-entropy loss gradient:
$\mathbb{E}_{E}\left[\frac{\partial\ell}{\partial\theta}\right]$ $=\mathbb{E}_{E}\left[\frac{1}{B}\sum_{\mu=1}^{B}\sum_{k=1}^{K}(c_{k}+e_{k}^{\mu})(p_{k}^{\mu}-\mathbf{1}_{y=k})\right]$
@@
By the orthogonality assumption on logit gradient means $c_{k}$, the norm of this quantity is $d\sigma_{c}^{2}\lVert\hat{Q}-Q\rVert$. In Figure 5b, we confirm this linear relationship on LeNet-5 shockingly well. Normally, randomly-initialized models are biased and have non-uniform average priors $\hat{Q}$, so that $\lVert\hat{Q}-Q\rVert>0$. In this case, we expect loss gradients computed on two potentially unrelated but balanced datasets to have considerable overlap as both should align with the non-zero expected gradient. Indeed, Figure 5a shows that gradients of LeNet-5 computed on CIFAR-10 and on random images drawn from a standard Gaussian have cosine similarity $>0.7$ when priors $Q$ and $\hat{Q}$ disagree, which is quite noteworthy for a 61,170-dimensional space. In contrast, for low-confident models that always predict uniform distribution (e.g., when $\alpha\ll 1$, see Figure 5a right), logit gradients cancel each other out when computed on a large balanced batch. Hence, rather counterintuitively, we derive that large positive curvature is associated with low-norm cross-entropy gradients for balanced datasets.
@@ 5 Connections to Optimization
We optimize $\alpha$-scaled homogeneous networks using vanilla full-batch gradient descent for 20,000 epochs. In particular, we use a fully-connected LeNet-300-100 with two hidden layers on FashionMNIST (Xiao et al., 2017) and a convolutional LeNet-5 with $4$ hidden layers on CIFAR-10 (LeCun et al., 1998; Krizhevsky, 2009), all implemented in PyTorch (Ansel et al., 2024). We set softmax temperature to $T=1$ to link the Goldilocks zone to the initialization norm. Recall from Section 3 that logit gradients of the $\alpha$-scaled network $f^{\prime}$ are $\alpha^{L-1}$ times the respective logit gradients of the unscaled model $f$, so we multiply a preset learning rate $\eta_{0}$ by $\alpha^{2-L}$ to ensure that updates are initially commensurate to the weights of $f^{\prime}$. In fact, this correction factor ensures that $f^{\prime}$ has exactly the same training dynamics as $f$ provided that logit gradients are combined using the same softmax output $p=p^{\prime}$ to compute the update vector as indicated in Equation 3. From this perspective, our experiments essentially constitute an ablation study examining the effects of extreme model (un)confidence, which corresponds to the exterior of the Goldilocks zone in our setup.
@@
Our findings are summarized in Figure 7. As long as the softmax ouptut remains uniform ($\alpha<1$), homogeneous models remain in the divergence regime characterized by an increase of the parameter norm in accordance with previous studies (Liu et al., 2023). Once the weights become sufficiently large for the predictions to become non-uniform, the network circles back to the Goldilocks zone and trains normally thereafter provided that the learning rate is approximately admissible, i.e., $\eta=\eta_{0}\alpha^{2-L}<2/\lVert H\rVert_{2}$ (Lewkowycz et al., 2020; Cohen et al., 2021; Wang et al., 2022). Hence, in Figure 8 (left) we observe a linear separation between trainable and divergent setups that extends well into the low-norm initialization region and way beyond the Goldilocks zone (see Appendix E for further discussion). When the learning rate is too high, however, the network breezes past the Goldilocks zone and either diverges to infinity or develops an increasing amount of zero logits (middle column in Figure 8). This happens as some negative weights grow much faster than the positive ones, promoting zero activations within the network by virtue of ReLU.
@@
The observations above reveal that homogeneous networks may not only converge when initialized outside the Goldilocks zone but also fail to learn when optimized well within its boundaries (Figure 8). The former occurs as the model circles back to the Goldilocks zone ($\alpha\ll 1$) or when it learns the lazy solution ($\alpha\gg 1$), and the latter—as it develops zero logits. Indeed, LeNet-5 does not perform above random for any $\alpha>1.8$ and any learning rate, whereas the corresponding Goldilocks zone spans $\alpha\in[1,18]$. Lastly, Figure 8 simply illustrates that phase transitions between different optimization regimes are not necessarily aligned with the Goldilocks zone, which is naturally highlighted in the right plots. Thus, we argue that positive curvature of the loss by itself is not a precise metric to evaluate the quality of initializations in terms of network trainability and generalization, and that further investigations are needed to clarify this connection.